Two hot dogs and one hamburger cost $3.39. Three hot dogs and three hamburgers cost $6.57. What is the cost of one hot dog?
Answer by Maths68(1474)

Let
hot dogs =x
hamburger = y
Two hot dogs and one hamburger cost $3.39
2x+y=3.39
y=3.39-2x........(1)
Three hot dogs and three hamburgers cost $6.57
3x+3y=6.57 ........(2)
Put the value of y from (1) to (2)
3x+3y=6.57 ........(2)
3x+3(3.39-2x)=6.57
3x+10.17-6x=6.57
-3x=6.57-10.17
-3x=-3.6
Divide by -3 both sides of above equation
-3x/-3=-3.6/-3
x=1.2
one hotdog cost = 1.2$
Put the value of x in (1)
y=3.39-2x
y=3.39-2(1.2)
y=3.39-2.4
y=0.99
Therefore
one hot dog =1.2$
one hamburger = 0.99$